OpenStudy (anonymous):

Write the equation of the hyperbola with foci at (1, 5) and (7, 5) and with vertices at (2, 5) and (6, 5). Answer x minus 4 quantity squared divided by 4 minus y minus 5 quantity squared divided by 5 equals 1 x plus 4 quantity squared divided by 4 minus y minus 5 quantity squared divided by 5 equals 1 x minus 4 quantity squared divided by 4 minus y plus 5 quantity squared divided by 5 equals 1 x minus 4 quantity squared divided by 5 minus y minus 5 quantity squared divided by 4 equals 1

OpenStudy (amistre64):

plug in your verts to see which option is true

OpenStudy (amistre64):

when y=5 .... what does your x values need to be?

OpenStudy (anonymous):

-4

OpenStudy (amistre64):

it hard to parse thru the options ... why are they worded and not notated?

OpenStudy (amistre64):

the center is (4,5) meaning that our x and y parts are going to be -4 and -5 that narrows it to the first and last options

OpenStudy (amistre64):

(x-4)^2/4^2 - (y-5)^2/5 = 1 (x-4)^2 has to equal 16 in this case (x-4)^2/ 5 - (y-5)^2/4 = 1 (x-4)^2 has to equal 5 in this case

OpenStudy (amistre64):

looks like a read that first one off a little .... its not 4^2, its just a 4 :)